OpenStudy (anonymous):

Evaluate a limit: limx→0 x −−−−−−−- √1−cosx

OpenStudy (anonymous):

x over sqrt of 1-cosx

OpenStudy (anonymous):

\[\LARGE \lim_{x \rightarrow 0} \frac{x}{\sqrt{1-\cos x}}\]

OpenStudy (anonymous):

\[\LARGE \lim_{x \rightarrow 0} \frac{x}{\sqrt{1-\cos x}} \times \frac{\sqrt{1+\cos x}}{\sqrt{1+\cos x}}=?\]

OpenStudy (anonymous):

I don't know the numerator. Could you teach me:)

OpenStudy (anonymous):

\[\LARGE \lim_{x \rightarrow 0} \frac{x}{\sqrt{1-\cos x}} \times \frac{\sqrt{1+\cos x}}{\sqrt{1+\cos x}}=\frac{x {\sqrt{1+\cos x }}}{\sqrt{1-\cos^2x}}\]

OpenStudy (anonymous):

\[\LARGE \frac{x {\sqrt{1+\cos x }}}{\sqrt{\sin^2x}}= \lim_{x \rightarrow 0} \frac{x \sqrt{1+\cos x}}{\sin x}\] and.. \[\LARGE \lim_{x \rightarrow 0} \frac{x}{\sin x}=1\]

OpenStudy (anonymous):

\[\Huge \lim_{x \rightarrow 0} \sqrt{1+ \cos x}\] Now you can put x=0

OpenStudy (anonymous):

did you get anything?

OpenStudy (anonymous):

Yes, I know what to do now, multiple by its conjugate again

OpenStudy (anonymous):

Thanks for your help:)

OpenStudy (anonymous):

u can just put x=0 \[\Huge \sqrt{1+\cos 0}\] cos0=?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

sqrt(1+1)=

OpenStudy (anonymous):

the answer should be zero since limit x/sinx = 0

OpenStudy (anonymous):

x/sinx=1..are you somewhat dizzy?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ x }{sinx }\]= \[\lim_{x \rightarrow 0}\frac{ 0 }{sin0 }\]=0/1=0?? If I am mistaken, please correct me

OpenStudy (anonymous):

That is a property,you don't have to put 0 in that! You just have to remember that when x tends to 0,the expression sinx/x or x/sinx tends to 1

OpenStudy (anonymous):

That is because for small x, sin x is approximately x, as can be seen in its power series expansion.